Required information Exercise 6-16 Working with a Segmented Income Statement; Break-Even Analysis (L06-4, LO6-5] [The followiExercise 6-16 Part 1 Required: 1-a. Compute the companywide break-even point in dollar sales. 1-b. Compute the break-even poiExercise 6-16 Part 1 Required: 1-a. Compute the companywide break-even point in dollar sales. 1-b. Compute the break-even poiExercise 6-16 Part 1 Required: 1-a. Compute the companywide break-even point in dollar sales. 1-b. Compute the break-even poiExercise 6-16 Part 2 2. By how much would the companys net operating income increase if Minneapolis increased its sales by $
Answer #1

Break even point = Fixed cost/Contribution margin ratio

1A.

Breakeven point in dollar sales = (126,000+63,000)/50%

= 378,000

.

1B.

Chicago= 78,000/70%

= 111,429

Minneapolis = 48,000/40%

= 120,000

.

1C. Greater

.

Net operating income increase = Increase in sales* contribution margin ratio

= 75,000*40%

= 30,000